OpenStudy (daenio):

Suppose spin a spinner with eight equal sectors, numbered 1 to 8. What is the probability of spinning a 3 or an 8?

Directrix (directrix):

There are 8 possible outcomes. Of these, two are favorable (the 3 and the 8). P(3 or 8) = 2/8 = 1/4

OpenStudy (daenio):

Hey man, thanks! Can you help me some other things?

Directrix (directrix):

I'll try. Post them in separate threads.

OpenStudy (daenio):

No problem, thanks!